Problem #14

A grain elevator has “short lived” components which require replacing every 5 years.  The cost is estimated at $60,000.00.  What amount should be budgeted annually to a 10% sinking fund to allow for recapture?
Answer #1

Annual amount to be deposited = 60000 * (A/F,10%,5) = 60000 * 0.163797 = 9827.82 ~ 9828